OpenStudy (anonymous):

John and David are running the 1-mile race. David enters the final quarter first and is running at a rate of 18 ft./sec. John enters the final quarter 3 seconds later but is running at a rate of 20ft./sec. Who will win the race? hint 5280ft.= 1mi.

OpenStudy (anonymous):

Well, a quarter of a mile is or can be given by (1/4)(5280), or 1320 feet

OpenStudy (anonymous):

We use a little formula here: D=R*T, which basically stand fro Distance =Rate*Time

OpenStudy (anonymous):

have you tried this problem on your own? what was your answer?

OpenStudy (anonymous):

yeah but i dont get it

OpenStudy (radar):

You understand who gets to the finish line first is the winner. Since David enters the final quarter mile (1320 ft) 3 seconds before John, he has a head start and has traveled 3 times his blistering speed of 18 ft. per second, or 54 feet, and has only 1320-54 feet or 1266 feet left when John enters the last quarter mi. Now the question is who gets there first? Will David cover his remaining 1266 feet at 18 ft/sec before John travels his 1320 ft at his rate of 20 ft/sec. You now have the data needed to solve. D=R*T Rearrange to solve for time getting T=D/R. Plug in the D and R values for David and then John and see which T is the lesser, that will be the winner. David T will be 1266/18 John T will be 1320/20 The moral of this story is he who gets to the last qtr first may not be the one who finishes first.